# {tocentry}

### v1

@@ 4.4 Energy Conservation
Figure 4: Energy conservation of the QM/MM-Multipole method. (A) The potential energy surface with a constant $R_{\mathrm{cut}}$ as a function of an MM water (shown as balls and sticks) distance from the QM cluster (shown as spheres) center. (B) Energy conservation in a QM-water-in-MM-water system. Linear fits to the energy time series are plotted as solid lines, and the text in the plots show the slopes of the fits. (C) Energy conservation in a chorismate mutase model.
Since we treat the electrostatics in regions I and II differently, the potential energy surface of our QM/MM-Multipole approach is not strictly smooth. In Figure 4A, we check the smoothness of the energy surface when one MM water is moved from region I to region II with a fixed $R_{\mathrm{cut}}=17$. In contrast to the large energy jumps observed for QM/MM-Ewald, we see smooth (on the scale of the plot) energies in the QM/MM-Multipole approach. The large jump of the QM/MM-Ewald energy right at $R=R_{\mathrm{cut}}$ occurs when the boundary cuts through the MM water molecule. This is because although the leading electrostatic interaction with a water molecule should only involve the water dipole, when the water molecule is cut, the electrostatics in QM/MM Ewald is separated into interactions with individual charges on the water molecule, some of which are treated exactly and some of which are treated in the QM monopole approximation. There is thus a large residual charge interaction term which is unphysical. The QM/MM Ewald energy still has a noticeable jump however, even when we compare energies just before and after a whole water molecule is moved to region II, when this charge interaction error is absent.
@@ 4.5 Example Application: Catalytic Kinetics in Chorismate Mutase
1HB/S+R90 $(5.1\pm 0.3)\times 10^{3}$ $(2.7\pm 0.4)\times 10^{-8}$ $(7\pm 1)\times 10^{-4}$
2HB/S+R90 $(1.1\pm 0.2)\times 10^{-4}$
2HB/S+R90+R7+E78 $(1.2\pm 0.2)\times 10^{7}$ $(5.8\pm 0.7)\times 10^{-4}$ $(1.7\pm 0.6)\times 10^{-1}$
@@
a The experimental catalytic rate at 300 K was computed from the fitted temperature-dependent rate $k_{\mathrm{cat}}=k_{B}T/h\exp{(-\Delta H/k_{B}T+\Delta S/k_{B})}$ with parameters $\Delta H=12.7\pm 0.4$ kcal/mol and $\Delta S=-9.1\pm 1.2$ cal/mol/K, taken from Ref.72, divided by the solvent kinetic isotope effect (KIE) measured at 303.15 K73. The error in this rate was estimated by the uncertainty propagation equation74 by assuming the errors of $\Delta H$ and $\Delta S$ are independent and no error in KIE. The normalization by KIE is for direct comparison to the theoretical rates obtained with classical nuclear dynamics.
@@
Table 2: Catalytic rate of chorismate mutase from Bacillus subtilis in s-1. Theoretical rate errors are reported as the square root of the rate estimator variance. Statistical analysis details are provided in the SI.
@@
Chorismate Mutase (CM) is a critical component in the primary and secondary metabolism of fungi, bacteria, and plants75. It catalyzes the Claisen rearrangement that transforms chorismate into prephenate (Figure 5A), an important precursor for the biosynthesis of aromatic amino acids76, 77, 75. The catalytic kinetics of this reaction were recently studied by Ray et al. via MD enhanced sampling simulations on a QM(PBE)/MM PES61. They computed the reaction rate using a variant of the flooding method78 called OPES flooding62, which adaptively adds bias potentials to the reactant state to accelerate barrier crossings that would otherwise be inaccessible to unbiased simulations. The computed QM/MM reaction rate was found to be somewhat higher than the experimentally measured catalytic rate, and this discrepancy was attributed by the authors to the neglect of the free energy cost for the substrate to adopt a “reactive” conformation.
@@
We now re-examine this reaction, following the simulation setup of Ray et al. as closely as possible, but further explore how the theoretical kinetic predictions depend on the MD sampling, the QM region definition, and the employed QM theory. In all our simulations, we used the same scheme to determine $R_{\mathrm{cut}}$ as in the energy conservation test. We verified our QM/MM scheme and implementation by reproducing the QM(PBE/DZVP)/MM result of CP2K by Ray et al. using our QM(PBE/6-31G**)/MM-Multipole MD. We found that the calculated rate ($5.1\times 10^{3}$ s-1) indeed overestimates the experimental rate of 16 s-1, while being of the same order of magnitude as the rate reported by Ray et al. (9433 s-1).
@@
We first consider how the rate depends on the level of quantum theory. To do so we used two more accurate hybrid density functionals: $\omega$B97X-3c and a refined version of it parametrized to reproduce the DLPNO-CCSD(T)/CBS reaction barriers for this system (see Computational Details). We kept the QM region definition unchanged and started from the same structure as Ray et al. (with appropriate equilibration using the corresponding functional). By changing the quantum description, we arrive at an opposite conclusion – the QM/MM reaction rate now significantly underestimates the experimental catalytic rate (the first row of Table 2). The qualitatively different conclusion is a result of the inaccuracy in the pure functional treatment of the QM region, with its substantial underestimate of the reaction barrier, as shown by the gas-phase energetic analysis (Figure 5H). Assuming that the experimental rate is exact and the refined $\omega$B97X-3c functional PES is almost exact, the deviation of the computed rate must be a result of either insufficient MD sampling (statistical error) or an inadequate QM region size (a source of systematic error).
@@
Figure 5: (A) The chorismate mutase catalyzed reaction. (B-D) Nudged elastic band optimized geometries of the reactant, the transition state and the product when Arg90 forms a single hydrogen bond with the substrate. (E-G) The optimized reactant, transition state and product geometries when Arg90 forms double hydrogen bonds with the substrate. (H) Gas-phase energies on NEB geometries for the 1HB conformation. (I) Gas-phase energies on NEB geometries for the 2HB conformation.
@@
As shown in Table 2, our calculated rate does not appear to suffer from a large statistical error, but in the simulations initiated from the conformation of Ref. 61, we never observed a double hydrogen bond between the catalytic Arg90 and the substrate, even though this is a known binding mode found in a CM crystal structure (PDB id: 1COM)79. Instead, Arg90 forms only one hydrogen bond with the substrate (Figure 5B), resembling the conformation captured in another crystal structure (PDB id: 2CHT)80. The main difference between the two binding modes is in the orientations of the $\epsilon$-nitrogen/hydrogen of Arg90. In 2CHT, a HB between the Arg90 terminal nitrogen and the substrate carboxylic oxygen is found with no other HB interactions between the two moieties. In 1COM, two HBs are found, one between the $\epsilon$-nitrogen and the chorismate carboxylic oxygen, and another between the terminal nitrogen and the chorismate ester oxygen. The transition from the 1HB conformation to the 2HB conformation requires a flip of Arg90, which is not accessible in our QM/MM equilibration.
@@
To investigate how the Arg90 conformation affects the reaction rate, we added biasing potentials to artificially accelerate this flip and then equilibrated the resulting 2HB structure after removing the bias (see Figure 5E for an example of the 2HB structures). We performed OPES flooding on this new structure to compute the rate. We found then that the calculated rate of the $\omega$B97X-3c functional was enhanced by four orders of magnitude (the second row of Table 2).
@@
The sensitivity of the rate to the local protein conformation underlines the challenges of converging the sampling, especially when one has no prior knowledge of the existence of alternative binding modes. Interestingly, the gas-phase single point energetics show a lower reaction barrier for the 1HB binding mode than for the 2HB binding mode (Figures 5H and 5I), contradicting the significantly faster reaction rate obtained from OPES flooding simulations with the 2HB conformation. This illustrates the importance of entropic contributions and the appropriate modeling of the chemical environment.
@@
It has been conjectured that the catalytic mechanism of CM involves an acid/base attack step due to the considerable measured solvent kinetic isotope effects (KIE)73 on the apparent catalytic rate, $k_{\mathrm{H_{2}O}}/k_{\mathrm{D_{2}O}}=2.23$. The dramatic dependency of the theoretical rate on the hydrogen-bonding pattern suggests, however, that the solvent KIE results from the hydrogen bond strength’s dependence on the quantum nature of the protons; further simulations with an explicit treatment of nuclear quantum effects (NQEs) should be carried out to explore this point.
We also tested the sensitivity of the rate to the definition of the QM region by further including Arg7 and Glu78 in the QM region. When switching these residues from an MM to a QM description, we found that the calculated rate changed only slightly, increasing by less than one order of magnitude. Mutagenesis studies on Arg7 and Glu78 reported substantially lower enzyme activities 81, 82. Within our study, the modest change in rates upon changing from a classical to quantum description of the residues suggests that they participate via classical electrostatics rather than by participating in charge transfer to the substrate or catalytic Arg90 residue.
@@
Our final theoretical rate with the refined $\omega$B97X-3c functional, the (larger) S+R90+R7+E78 QM region definition, and the double hydrogen-bonded Arg90 conformation, shows a reasonable agreement with the estimated experimental rate. At the same time, our simulations illustrate the difficulties of obtaining a fully converged result. Improving this agreement will require improving the errors in the theoretical treatment, from the imperfect electronic structure, size of QM region, sampling, and nuclear quantum effects. Within DFT treatments, the quality of the functional remains perhaps the largest source of uncertainty, with almost 11 orders of magnitude difference in rate between different functionals.
@@
From the PBE data, where the rate with the 2HB Arg90 conformation and the larger QM region is much faster than the experimental rate, one might conclude that the Claisen rearrangement step is not rate-limiting for the whole catalytic cycle, and perhaps the substrate leaving step is instead rate-limiting, as has been hypothesized in some earlier experiments72. However, our results using the improved PES provided by the refined $\omega$B97X-3C functional suggest the chemical transformation step is rate-limiting regardless of the Arg90 conformation or the QM region definition. This aligns with the proposal that electrostatic transition state stabilization is fundamental for CM catalysis83, and consistent with data that finds impaired enzyme activity after mutating Arg90 into a neutral citrulline, an arginine analog that is still able to preorganize the chorismate into its “reactive” conformation.
@@ 5 Conclusions
In this work, we described a QM/MM implementation with an improved treatment of the periodic QM/MM electrostatics and which leverages recent advances in GPU DFT implementations to enable the practical use of advanced hybrid density functionals. Our benchmarks verify its numerical stability and illustrate the possibility to control and converge the error associated with the electrostatics.
While this work has not focused on achieving a computationally optimal implementation, the open-source availability of our implementation and already competitive level of performance achieved forms a foundation for broader community efforts to build an even faster QM/MM code.

### v2

@@ 4.4 Energy Conservation
Figure 4: Energy conservation of the QM/MM-Multipole method. (A) The potential energy surface with a constant region I $R_{\mathrm{cut}}$ as a function of an MM water (shown as balls and sticks) distance from the QM cluster (shown as spheres) center. (B) Energy conservation in a QM-water-in-MM-water system. Linear fits to the energy time series are plotted as solid lines, and the text in the plots shows the slopes of the fits. (C) Energy conservation in a chorismate mutase model.
Since we treat the electrostatics in regions I and II differently, the potential energy surface of our QM/MM-Multipole approach is not strictly smooth. In Figure 4A, we check the smoothness of the energy surface when one MM water is moved from region I to region II with a fixed $R_{\mathrm{cut}}=17$. In contrast to the large energy jumps observed for QM/MM-Ewald, we see smooth (on the scale of the plot) energies in the QM/MM-Multipole approach. The large jump of the QM/MM-Ewald energy right at $R=R_{\mathrm{cut}}$ occurs when the boundary cuts through the MM water molecule. This is because although the leading electrostatic interaction with a water molecule should only involve the water dipole, when the water molecule is cut, the electrostatics in QM/MM Ewald is separated into interactions with individual charges on the water molecule, some of which are treated exactly and some of which are treated in the QM monopole approximation. There is thus a large residual charge interaction term which is unphysical. The QM/MM Ewald energy still has a noticeable jump, however, even when we compare energies just before and after a whole water molecule is moved to region II, when this charge interaction error is absent.
@@ 4.5 Example Application: Catalytic Kinetics in Chorismate Mutase
1HB/S+R90 $(5.1\pm 0.3)\times 10^{3}$ $(2.7\pm 0.4)\times 10^{-8}$ $(4.0\pm 0.7)\times 10^{-3}$
2HB/S+R90 $(9.2\pm 0.5)\times 10^{5}$ $(1.1\pm 0.2)\times 10^{-4}$ $(4\pm 2)\times 10^{-2}$
2HB/S+R90+R7+E78 $(1.2\pm 0.2)\times 10^{7}$ $(5.8\pm 0.7)\times 10^{-4}$ $1.1\pm 0.2$
@@
a The experimental catalytic rate constant at 300 K was computed from the fitted temperature-dependent rate constant $k_{\mathrm{cat}}=k_{B}T/h\exp{(-\Delta H/k_{B}T+\Delta S/k_{B})}$ with parameters $\Delta H=12.7\pm 0.4$ kcal/mol and $\Delta S=-9.1\pm 1.2$ cal/mol/K, taken from Ref.92, divided by the solvent kinetic isotope effect (KIE) measured at 303.15 K93. The error in this rate constant was estimated by the uncertainty propagation equation94 by assuming the errors of $\Delta H$ and $\Delta S$ are independent and no error in KIE. The normalization by KIE is for direct comparison to the theoretical rate constants obtained with classical nuclear dynamics.
@@
Table 2: Catalytic rate constant of chorismate mutase from Bacillus subtilis in s-1. Theoretical rate constant errors are reported as the square root of the rate constant estimator variance. Statistical analysis details are provided in the SI.
@@
Chorismate Mutase (CM) is a critical component in the primary and secondary metabolism of fungi, bacteria, and plants95. It catalyzes the Claisen rearrangement that transforms chorismate into prephenate (Figure 5A), an important precursor for the biosynthesis of aromatic amino acids96, 97, 95. The catalytic kinetics of this reaction were recently studied by Ray et al. via MD enhanced sampling simulations on a QM(PBE)/MM PES79. They computed the reaction rate using a variant of the flooding method98 called OPES flooding80, which adaptively adds bias potentials to the reactant state to accelerate barrier crossings that would otherwise be inaccessible to unbiased simulations. The computed QM/MM reaction rate constant was found to be somewhat higher than the experimentally measured catalytic rate constant, and this discrepancy was attributed by the authors to the neglect of the free energy cost for the substrate to adopt a “reactive” conformation.
@@
We now re-examine this reaction, following the simulation setup of Ray et al. as closely as possible, but further explore how the theoretical kinetic predictions depend on the MD sampling, the QM region definition, and the employed QM theory. In all our simulations, we used the same scheme to determine $R_{\mathrm{cut}}$ as in the energy conservation test. We verified our QM/MM scheme and implementation by reproducing the QM(PBE/DZVP)/MM result of CP2K by Ray et al. using our QM(PBE/6-31G**)/MM-Multipole MD. We found that the calculated rate constant ($5.1\times 10^{3}$ s-1) indeed overestimates the experimental one of 16 s-1 while being of the same order of magnitude as the value reported by Ray et al. (9433 s-1).
@@
We first consider how the rate constant depends on the level of quantum theory. To do so we used two more accurate hybrid density functionals: $\omega$B97X-3c and a refined version of it parametrized to reproduce the LNO-CCSD(T)/CBS reaction barriers for this system (see Computational Details). We kept the QM region definition unchanged and started from the same structure as Ray et al. (with appropriate equilibration using the corresponding functional). By changing the quantum description, we arrive at an opposite conclusion – the QM/MM reaction rate constant now significantly underestimates the experimental catalytic rate constant (the first row of Table 2). The qualitatively different conclusion is a result of the inaccuracy in the pure functional treatment of the QM region, with its substantial underestimate of the reaction barrier, as shown by the gas-phase energetic analysis (Figure 5H). Assuming that the estimated experimental rate constant is accurate and the refined $\omega$B97X-3c functional PES is almost exact, the deviation of the computed rate constant must be a result of either insufficient MD sampling (statistical error) or an inadequate QM region size (a source of systematic error).
@@
Figure 5: (A) The chorismate mutase catalyzed reaction. (B-D) Nudged elastic band optimized geometries of the reactant, the transition state, and the product when Arg90 forms a single hydrogen bond with the substrate. (E-G) The optimized reactant, transition state, and product geometries when Arg90 forms double hydrogen bonds with the substrate. Energies are aligned at the reactant state. (H) Gas-phase energies on NEB geometries for the 1HB conformation. (I) Gas-phase energies on NEB geometries for the 2HB conformation.
@@
As shown in Table 2, our calculated rate constant does not appear to suffer from a large statistical error, but in the simulations initiated from the conformation of Ref. 79, we never observed a double hydrogen bond between the catalytic Arg90 and the substrate, even though this is a known binding mode found in a CM crystal structure (PDB id: 1COM)99. Instead, Arg90 forms only one hydrogen bond with the substrate (Figure 5B), resembling the conformation captured in another crystal structure (PDB id: 2CHT)100 that was used to initiate the simulations of Ray et al. The main difference between the two binding modes is in the orientations of the $\epsilon$-nitrogen/hydrogen of Arg90. In 2CHT, an HB between the Arg90 terminal nitrogen and the substrate carboxylic oxygen is found with no other HB interactions between the two moieties. In 1COM, two HBs are found, one between the $\epsilon$-nitrogen and the chorismate carboxylic oxygen, and another between the terminal nitrogen and the chorismate ester oxygen. The transition from the 1HB conformation to the 2HB conformation requires a flip of Arg90, which is not accessible in our QM/MM equilibration.
@@
To investigate how the Arg90 conformation affects the reaction rate, we added biasing potentials to artificially accelerate this flip and then equilibrated the resulting 2HB structure after removing the bias (see Figure 5E for an example of the 2HB structures). We performed OPES flooding on this new structure to compute the rate constant. We found then that the calculated rate constants were enhanced regardless of the functional choice, with the $\omega$B97X-3c one even increasing by four orders of magnitude (the second row of Table 2).
@@
The sensitivity of the rate to the local protein conformation underlines the challenges of converging the sampling, especially when one has no prior knowledge of the existence of alternative binding modes. Interestingly, the gas-phase single point energetics show a lower reaction barrier for the 1HB binding mode than for the 2HB binding mode (Figures 5H and 5I), contradicting the faster reaction rate constants obtained from OPES flooding simulations with the 2HB conformation. This illustrates the importance of entropic contributions and the appropriate modeling of the chemical environment.
@@
It has been conjectured that the catalytic mechanism of CM involves an acid/base attack step due to the considerable measured solvent kinetic isotope effects (KIE)93 on the apparent catalytic rate constant, $k_{\mathrm{H_{2}O}}/k_{\mathrm{D_{2}O}}=2.23$. The sensitivity of the theoretical rate constant on the hydrogen-bonding pattern suggests, however, that the solvent KIE results from the hydrogen bond strength’s dependence on the quantum nature of the protons; further simulations with an explicit treatment of nuclear quantum effects (NQEs) should be carried out to explore this point.
We also tested the sensitivity of the rate constant to the definition of the QM region by further including Arg7 and Glu78 in the QM region. When switching these residues from an MM to a QM description, we found all the calculated rate constants further increase (the third row of Table 2), consistent with substantially lower enzyme activities in mutagenesis analyses of Arg7 and Glu78 101, 102. The QM treatment allows a proton on Arg90 to be shared with Glu78 in PBE/6-31G** simulations, and even occasionally, for a proton transfer (PT) from Arg90 to Glu78 (1 in 11 simulations of $\omega$B97X-3c and 4 in 11 simulations of the refined $\omega$B97X-3c). However, the difference in the $\ln{k}$ of the refined $\omega$B97X-3c with 2HB/S+R90+R7+E78 is not statistically significant between the runs where PT happens and the ones where it does not (the $p$-value from a Welch’s $t$-test is 0.36 assuming equal means as the null hypothesis). This lack of statistical significance suggests that the enhanced rate cannot be directly attributed to the occurrence of PT alone. Instead, the observed rate increase may be explained by other factors, including the potential charge transfer and the differences in the electrostatic representation between QM and classical MM charges.
@@
Our final theoretical rate constant with the refined $\omega$B97X-3c functional, the (larger) S+R90+R7+E78 QM region definition, and the double hydrogen-bonded Arg90 conformation, shows a reasonable agreement with the estimated experimental one. At the same time, our simulations illustrate the difficulties of obtaining a fully converged result. Improving this agreement will require improving the errors in the theoretical treatment, from the imperfect electronic structure, size of QM region, sampling, nuclear quantum effects, and probably the rate theory as well (the correct estimate of rates depends on a careful choice of OPES flooding parameters according to Ref. 80). Within DFT treatments, the quality of the functional remains perhaps the largest source of uncertainty, with almost 11 orders of magnitude difference in rate constant between different functionals.
@@
From the PBE data, where the rate constant with the 2HB Arg90 conformation and the larger QM region is much faster than the experimental value, one might conclude that the Claisen rearrangement step is not rate-limiting for the whole catalytic cycle, and perhaps the substrate leaving step is instead rate-limiting, as has been hypothesized in some earlier experiments92. However, our results using the improved PES provided by the refined $\omega$B97X-3C functional suggest the chemical transformation step is rate-limiting regardless of the Arg90 conformation or the QM region definition. This aligns with the proposal that electrostatic transition state stabilization is fundamental for CM catalysis103, and consistent with data that finds impaired enzyme activity after mutating Arg90 into a neutral citrulline, an arginine analog that is still able to preorganize the chorismate into its “reactive” conformation.
@@ 5 Conclusions
In this work, we described a QM/MM implementation with an improved treatment of the periodic QM/MM electrostatics which leverages recent advances in GPU DFT algorithms to enable the practical use of advanced hybrid density functionals. Our benchmarks verify its numerical stability and illustrate the possibility of controlling and converging the error associated with the electrostatics.
While this work has not focused on achieving a computationally optimal implementation, the open-source availability of our code and the already competitive level of performance achieved form a foundation for broader community efforts to build an even faster QM/MM code.
@@
Supporting Information available. Details on multipolar Ewald and pseudo-bond parameterization, additional computational details on chorismate mutase, rate calculations, and statistical error analysis.